OpenStudy (anonymous):

why does n!/(n+1)!=1/(n+1)?

OpenStudy (anonymous):

it doesnt?

OpenStudy (anonymous):

My calculus book says tit does :(

jimthompson5910 (jim_thompson5910):

n!/(n+1)!=1/(n+1) n!/[(n+1)n!]=1/(n+1) 1/(n+1)=1/(n+1)

OpenStudy (anonymous):

So (n+1)!=(n+1)n!?

OpenStudy (anonymous):

I cant find any "rules" of factorials online

jimthompson5910 (jim_thompson5910):

yes, this is because n! = n*(n-1)*(n-2)...3*2*1 so (n+1)! = (n+1)*n*(n-1)*(n-2)...3*2*1 (n+1)! = (n+1)*[ n*(n-1)*(n-2)...3*2*1 ] (n+1)! = (n+1)*[ n! ] (n+1)! = (n+1)*n!